Newton law of universal gravitational is represented by F=GMm/r2 where G is constant
The two Mm are mass
r=radius.Determine the units of G

Grade:12th pass

2 Answers

Rani Goswami
2 years ago
Hiee:)
F= kg m/s^2 
Kg m/s^2 = Kg^2 m^-2 G 
Solving this G = kg^-1 m^3 s^-2
 
In dimensions it will be M^-1 L^3 T ^-2
Peace😊
Alpha Array
2 years ago
As we know that
F = GMm/r²
or,  G = Fr²/Mm
Now, units of G,
[Force→N ; r²→m² ; Mm→Kg²]
G = Nm²/Kg²
or, it can also be written as,
G = Mm²Kg-²